TIFFCMP(1) TIFFCMP(1) NAME tiffcmp−comparetwoTIFFfiles SYNOPSIS tiffcmp[options]file1.tiffile2.tif DESCRIPTION TiffcmpcomparesthetagsanddataintwofilescreatedaccordingtotheTaggedImageFileFormat,Revision6.0.Theschemesusedforcompressingdataineachfileareimmaterialwhendata arecompared−dataarecomparedonascanline-by-scanlinebasisafterdecompression.Most directorytagsarechecked;notableexceptionsare:GrayResponseCurve,ColorResponseCurve,and ColorMaptags.DatawillnotbecomparedifanyoftheBitsPerSample,SamplesPerPixel,or ImageWidthvaluesarenotequal.Bydefault,tiffcmpwillterminateifitencountersanydifference. OPTIONS −lListeachbyteofimagedatathatdiffersbetweenthefiles. −tIgnoreanydifferencesindirectorytags. BUGS Tagsthatarenotrecognizedbythelibraryarenotcompared;theymayalsogeneratespurious diagnostics. SEEALSO pal2rgb(1),tiffinfo(1),tiffcp(1),tiffmedian(1),libtiff(3) September26,1994 1